If the equation of a line is y=4/3x+8 what is the equation of a line six away and parallel to the first line? It has to do something with being six away based off the length of the perpendicular.
Answer by lwsshak3(11628)

If the equation of a line is y=4/3x+8 what is the equation of a line six away and parallel to the first line? It has to do something with being six away based off the length of the perpendicular.
**
y=4/3x+8
The problem is to determine how much to drop given equation so that the perpendicular distance from a parallel line is 6 units.
let x=distance y must be dropped
You can work with a right triangle in which this perpendicular distance is the opposite side of a right triangle whose tan of the opposite angle is 3/4, the slope of given straight line equation.
The sin of this angle is equal to 3/5.
sin of angle=3/5=6/x
x=10
Given equation should be dropped vertically 10 units
Ans:
Equation of a line six away and parallel to the first line: y=4x/3-2